OpenStudy (anonymous):

How can I find all the zeros of the the function f(x) = 9x^3 - 7x -2?

OpenStudy (anonymous):

Factor it.

OpenStudy (anonymous):

Or graph it. This does not look nice to factor.

OpenStudy (anonymous):

Oh I see.. \((x-1)(3x+1)(3x+2)\)

OpenStudy (anonymous):

Isnt there a step in between,, like : 9x^2-6x-3=0

OpenStudy (anonymous):

3x^2-2x-1=0

OpenStudy (anonymous):

How did you propose graphing it without the factors, anyway, polpak?

OpenStudy (anonymous):

Let me start over... shoudnt it be 9x(3)-7x-2 the zeros will be 1, -2/3, -1/3

OpenStudy (anonymous):

that's what calculators are for ;p

OpenStudy (anonymous):

And MDot, the first step is noting that f(1) = 0 (by inspection), so (x-1) is a factor. Then youjust go normally.

OpenStudy (anonymous):

And get to the above thing, probably with (x-1)(quad) which becomes the thing above

OpenStudy (anonymous):

ok so, (0)= -2 (+/- 1), (+/- 2)

OpenStudy (anonymous):

and q is a factor of a(n)= 9 right?

OpenStudy (anonymous):

I have no idea what you were talking about, I was just giving you the steps polpak (presumably) went through, when you asked wasn't there something in the middle. your zeros 1, -2/3, -1/3 before were right...

OpenStudy (anonymous):

alright, well i guess ill keep working it